# Building-block flow model for computational fluids

### v1

@@ 3.1 Opportunities of BFM for CFD
A distinctive advantage of the BFM, which might contribute to addressing the aforementioned challenges compared to previous models, is its ability to accommodate various flow regimes through the BBF database. The current version of BFM includes separated flow and the effects of zero, adverse, and favorable mean pressure gradients. However, the process of incorporating additional flow physics is scalable by expanding the building block database, a capability absent in current SGS/wall models. Future versions of BFM could involve BBFs accounting for laminar flow, compressibility effects, shock waves, wall roughness, and the laminar-to-turbulent transition. Efforts to include these new physics are already underway in our group.
@@
3.2 Conclusions
We have introduced the Building-Block Flow Model (BFM) for large-eddy simulation. The model is designed to address challenges encountered in CFD within the aerospace industry, specifically the demand for accurate and robust solutions at an affordable computational cost. The core assumption of the BFM is that the subgrid-scale physics of complex flows can be effectively represented by the physics of simpler canonical flows. Implemented using three feedforward ANNs for GPU architectures, the BFM is applicable to arbitrarily complex geometries. Unlike previous models, the training data for the BFM is derived from controlled WMLES with an ‘exact’ model for the mean quantities of interest. This approach ensures consistency with the numerical discretization and grid structure of the solver. We have shown that, at the resolutions considered here, the BFM offers improvements in predictions over conventional SGS and wall models in both simple and complex scenarios.
Truly revolutionary improvements in WMLES will encompass advancements in numerics, grid generation, and wall/SGS modeling. The BFM addresses all these aspects by devising SGS/wall models consistent with the numerics and the grid. The modularity of the BFM opens up new opportunities for developing SGS/wall models that have broad applicability across different scenarios. In essence, the BFM offers a single model that can accurately represent various flow physics, eliminating the need for multiple specialized models for specific flow types. To enhance model performance, we will continue the training of future versions of the BFM with additional building blocks to account for a richer collection of flow physics.
Acknowledgments
This work was supported by the National Science Foundation under grant #2317254 and by an Early Career Faculty grant from NASA’s Space Technology Research Grants Program (grant #80NSSC23K1498). S. C was supported by The Boeing Company. G. A. was partially supported by the NNSA Predictive Science Academic Alliance Program (PSAAP; grant DE-NA0003993). The authors acknowledge the support of the CTR Summer Program 2022 at Stanford University. The authors also acknowledge the Massachusetts Institute of Technology, SuperCloud, and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported here.
Author contributions
Competing interests
The authors declare no competing interests.
Appendix A Methods
A.1 Numerical solver and conventional models
@@ A.2 BFM formulation
The ANNs are fully-connected feedforward networks. The first ANN, tasked with predicting the wall-shear stress, consists of 6 layers with 40 neurons per layer. The second near-wall ANN consists of 10 layers with 12 neurons per each layer; and the outer-layer ANN consists of 10 layers with 16 neurons per layer. The non-dimensionalization of the input and output features is attained by using parameters that are local in both time and space to guarantee the applicability of the model to complex geometries. For the ANN responsible for predicting the wall-shear stress, the input and output quantities are non-dimensionalized using viscous scaling: $\nu$ and $\Delta$. For the ANNs predicting the eddy viscosity, the input and output variables are non-dimensionalized using semi-viscous scaling: $(\nu\sqrt{\mathbf{\tilde{S}}:\mathbf{\tilde{S}}})^{1/2}$ and $\Delta$.
A.3 Training data preparation
The mean velocity profiles and wall-shear stress are extracted from DNS data of the BBFs. The turbulent channel flows are used to model the regime where turbulence is fully developed without significant mean-pressure gradient effects. Separation, adverse and favorable mean pressure gradient effects are learned from turbulent Poiseuille-Couette (TPC) flows. In these cases, the top wall moves at a constant speed ($U_{t}$) in the streamwise direction, and a mean pressure gradient is applied in the direction opposed to the top wall velocity. The pressure gradient ranges from mild to strong, so that the flow “separates” (i.e., zero wall shear stress) on the bottom wall. Favorable mean pressure gradient effects are obtained from the upper wall of the TPC cases.

### v2

@@ Discussion
A distinctive advantage of the BFM, which might contribute to addressing the aforementioned challenges compared to previous models, is its ability to accommodate various flow regimes through the BBF database. The current version of BFM includes separated flow and the effects of zero, adverse, and favorable mean pressure gradients. However, the process of incorporating additional flow physics is scalable by expanding the building block database, a capability absent in current SGS/wall models. The impact of excluding BBFs from the BFM training set is illustrated in the Methods section for the Gaussian Bump. Future versions of BFM could involve BBFs accounting for laminar flow, compressibility effects, shock waves, wall roughness, and the laminar-to-turbulent transition. Efforts to include these new physics are already underway in our group.
@@
Conclusions
We have introduced the Building-Block Flow Model (BFM) for large-eddy simulation. The model is designed to address challenges encountered in CFD within the aerospace industry, specifically the demand for accurate and robust solutions at an affordable computational cost. The core assumption of the BFM is that the subgrid-scale physics of complex flows can be effectively represented by the physics of simpler canonical flows. Implemented using three feedforward ANNs for GPU architectures, the BFM is applicable to arbitrarily complex geometries. Unlike previous models, the training data for the BFM is derived from controlled WMLES with an ‘exact’ model for the mean quantities of interest. This approach ensures consistency with the numerical discretization and grid structure of the solver. We have shown that, at the grid resolutions considered here, the BFM matches or improves the predictions from conventional SGS and wall models in both simple and complex scenarios.
Truly revolutionary improvements in WMLES will encompass advancements in numerics, grid generation, and wall/SGS modeling. The BFM addresses all these aspects by devising SGS/wall models consistent with the numerics and the grid. The modularity of the BFM opens up new opportunities for developing SGS/wall models that have broad applicability across different scenarios. In essence, the BFM offers a single model that can accurately represent various flow physics, eliminating the need for multiple specialized models for specific flow types. To enhance model performance, we will continue the training of future versions of the BFM with additional building blocks to account for a richer collection of flow physics and grid resolutions to improve its robustness.
Methods
Numerical solver and conventional models
@@ Methods
The ANNs are fully-connected feedforward networks. The first ANN, tasked with predicting the wall-shear stress, consists of 6 layers with 40 neurons per layer. The second near-wall ANN consists of 10 layers with 12 neurons per each layer; and the outer-layer ANN consists of 10 layers with 16 neurons per layer. The rationale for dividing the model into three ANNs is twofold: 1) task division (wall model versus SGS stress model), and 2) compatibility constraints (between the wall model and the near-wall SGS stress model). For the former, the initial ANN is responsible for predicting the wall shear stress (acting as a wall model), while the other two ANNs predict the eddy viscosity $\nu_{t}$. The second point relates to the ANN calculating $\nu_{t}$ at the control volumes adjacent to the wall. This ANN is not only tasked with obtaining $\nu_{t}$ for accurate mean flow predictions but also with ensuring compatibility with the wall model predictions, as the latter are tightly coupled with $\nu_{t}$ just above the boundary. A consequence of this coupling is that the output from the near-wall ANN scales differently from that of the ANN dedicated to the rest of the flow. Although it is feasible to train one single ANN for $\nu_{t}$, it was found that a better-performing model could be trained by separating the ANN into two.
The non-dimensionalization of the input and output features is attained by using parameters that are local in both time and space to guarantee the applicability of the model to complex geometries[58]. For the ANN responsible for predicting the wall-shear stress, the input and output quantities are non-dimensionalized using viscous scaling: $\nu$ and $\Delta$. For the ANNs predicting the eddy viscosity, the input and output variables are non-dimensionalized using semi-viscous scaling: $(\nu\sqrt{\mathbf{\tilde{S}}:\mathbf{\tilde{S}}})^{1/2}$ and $\Delta$, where the former represents a local velocity scale akin to the wall friction velocity $u_{\tau}$ typically used in wall-bounded flows, and the use of $\Delta$ as a length scale allows for accommodating the effect of different grid resolutions.
Training data preparation
The mean velocity profiles and wall-shear stress are extracted from DNS data of the BBFs. The turbulent channel flows are used to model the regime where turbulence is fully developed without significant mean-pressure gradient effects. The core region of the turbulent channel, where mean shear effects are weak, serves to model the SGS physics of isotropic turbulence. Separation, adverse and favorable mean pressure gradient effects are learned from turbulent Poiseuille-Couette (TPC) flows. In these cases, the top wall moves at a constant speed ($U_{t}$) in the streamwise direction, and a mean pressure gradient is applied in the direction opposed to the top wall velocity. The pressure gradient ranges from mild to strong, so that the flow “separates” (i.e., zero wall shear stress) on the bottom wall. Favorable mean pressure gradient effects are obtained from the upper wall of the TPC cases.
@@
Other approaches have been proposed in the literature to enforce compatibility between models and numerical schemes, most of them in the context of the RANS equations [47, 99]. Advances in model-consistent approaches for LES are more limited due to the chaotic nature of the system and the inherent inconsistency between the filtered Navier-Stokes and the LES equations. Noteworthy examples include reinforcement learning [65, 100, 101], filter-consistent formulations [78, 79], and optimal LES formulations [102, 103, 74].
Gaussian bump: details of the computational set-up
The computational domain and set-up is as in [90]. The domain is a rectangular prism that extends from $-L$ to $1.5L$ in the streamwise direction (with respect to the bump apex), ±$\pm 0.5L$ in the spanwise direction, and from $0L$ to $0.5L$ in the vertical direction. The lateral and top boundaries are free-slip, a constant uniform inflow is imposed at the inlet, and the non-reflecting characteristic boundary condition with constant pressure is applied at the outlet.
For the results presented in Fig. 6, the grid has three levels of isotropic refinement. Each refinement level has roughly 10 control volumes along the wall-normal direction and the average size of each level is twice the size of the previous level (see Fig. 6). The grid size of the layer closer to the wall is the smallest with $\Delta_{\min}\approx 0.026h$, where $h=0.0838L$ is the bump height. The number of control volumes is 8.7 million. This resolution yields of the order of 5 points per boundary layer at the bump apex. The simulations have been performed with a varying time step to ensure that the Courant-Friedrichs-Lewy number is less than 2.
Effect of the grid size
[…]
Comparison with RANS and WMLES
Figure 10(C,D) displays the $C_{p}$ and $C_{f}$ values obtained using DSM-EQ and VRE-EQ on grids with 452 million control volumes ($\Delta_{\min}\approx 0.003h$), as reported in ref. [90]. The figure also includes results from RANS simulations using the Spalart-Allmaras (S-A) model[104] and the (S-A)-RC-QCR2000[105], performed by Iyer & Malik[89] on a grid with 21 million cells (simulating only the half span), which is effectively five times the grid size used by BFM. While all simulations show convergence upstream of the bump, VRE-EQ and S-A fail to capture the flow separation; the (S-A)-RC-QCR2000 predicts separation but inaccurately in terms of $C_{p}$ within the bubble and $C_{f}$ downstream. Only DSM-EQ with 452 million control volumes follows closely with the experimental results. Remarkably, comparable accuracy was achieved by BFM on a grid with fewer than 9 million control volumes, which contains 52 times fewer control volumes.
Effect of the building-blocks flows on the BFM accuracy
We demonstrate that excluding BBFs modeling favorable/adverse pressure gradients from the training set significantly diminishes the performance of BFM on the Gaussian bump, as shown in Figures 10(E,F). The grid resolution matches that of Figure 6. The green line represents a version of BFM with the SGS trained solely on BBFs derived from turbulent channel flows, without considering BBFs that account for the effects of favorable/adverse pressure gradients. In the absence of these specific BBFs, the model fails to accurately capture flow separation. These results illustrate how the BFM framework can systematically incorporate additional flow physics by integrating new BBFs, a capability not explicitly demonstrated by previous ANN-based SGS models.
Figure 10: (A,C,E) Wall pressure and (B,D,F) wall friction coefficients along the $y/L=0$ plane. In (A,B), the results are obtained in a grid with $\Delta_{\min}\approx 0.015h$. Legend as in Fig. 6. In (C,D), the lines correspond to DSM-EQ (red) and VRE-EQ (yellow) in a grid with $\Delta_{\min}\approx 0.003h$ ($\approx$450 million control volumes)[90]. The symbols are S-A (squares), (S-A)-RC-QRC2000 (triangles). The dashed blue line is BFM from Fig. 6. In (E,F), the lines correspond to BFM (blue) and BFM trained without the building blocks accounting for adverse and favorable mean pressure gradient effects (green). White circles are experimental values[86].
CRM-HL: details of the computational set-up
We follow the computational setup from [106]. A semi-span aircraft geometry is simulated in a hemi-sphere of radius $1000c_{a}$, where $c_{a}$ is the mean aerodynamic chord. The symmetry plane is treated with free-slip and no penetration boundary conditions. A uniform plug flow is used at the front half of the hemisphere. A non-reflecting boundary condition with specified freestream pressure is imposed at the rear half of the hemisphere. Several grid refinements are considered, as illustrated in Fig. 2, being the coarsest and the smallest grid elements $\Delta_{\max}\approx c_{a}$ and $\Delta_{\min}\approx 2\times 10^{-3}c_{a}$, respectively. This leads to a total number of grid points is 30 million and the number of grid points per boundary layer thickness ranges from zero at the leading edge of the wing to twenty at the trailing edge. The reader is referred to [106] for more details about the numerical setup and gridding strategy.
Juncture flow: details of the computational set-up
We follow the computational setup from [26] The aircraft is centered in a rectangular prism whose sides are about 5 times the fuselage length, $L$, in the three directions. A uniform plug flow is imposed at the bottom and front boundaries; a non-reflecting boundary condition is used at the top and rear boundaries; and the lateral sides are modeled as free-slip. The mesh is constructed using a Voronoi diagram with a background grid size of $\Delta\approx 0.04L$, and several layers of refinement around the aircraft that leads to minimum grid size over the aircraft’s surface of $\Delta_{\min}\approx 4\times 10^{-4}L$. The reader is referred to ref.[26] for more details about the numerical setup and gridding strategy.
Acknowledgments
This work was supported by the National Science Foundation under grant #2317254 and by an Early Career Faculty grant from NASA’s Space Technology Research Grants Program (grant #80NSSC23K1498). S. C was supported by The Boeing Company. G. A. was partially supported by the NNSA Predictive Science Academic Alliance Program (PSAAP; grant DE-NA0003993). The authors acknowledge the support of the CTR Summer Program 2022 at Stanford University. The authors also acknowledge the Massachusetts Institute of Technology, SuperCloud, and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the research results reported here. Finally, the authors want to thank Rahul Agrawal for sharing the data of the Gaussian bump, and G.A. wishes to thank Prof. Manuel García-Villalba for insightful discussions.
Author contributions
Competing interests
The authors declare no competing interests.